# A short note on real algebraic maps which are topologically special generic maps

### 17 Dec 2023 v1

@@ Abstract.
Morse functions with exactly two singular points on homotopy spheres and canonical projections of spheres are generalized as special generic maps. A special generic map is, rounghly, a smooth map represented as the composition of a smooth surjection onto a manifold whose preimages are diffeomorphic to a unit sphere in the interior of the manifold and single point sets on the boundary with a smooth immersion of codimension 00.
@@
In theory of Morse functions, related singularity theory and applications to geometry of manifolds, Morse functions with exactly two singular points on spheres (in Reeb’s theorem) are of most fundamental functions and maps. Such maps are generalized to special generic maps. Special generic maps are smooth maps which are locally represented as projections or singular points of the form $(x_{1},\cdots x_{m})\mapsto(x_{1},\cdots x_{n-1},{\Sigma}_{j=1}^{m-n+1}{x_{j}}^{2})$. [4, 7, 27] are related pioneering studies. For theory of Morse functions, see [25] for a pioneering study and [21] for systematic exposition on fundamental notions and arguments.
@@
For example, [19, 24, 32] shows explicit cases and some fundamental properties of real algebraic functions in the theory of Lie groups and symmetric spaces. (Unit) spheres are seen as simplest symmetric spaces. We are interested in another generalizations of projections of spheres. We are interested in constructing real algebraic maps topologically special generic maps. The author has succeeded in such construction in [10], followed by [14] by the author himself. We present results of a similar type which are also new.
@@
Let $X$ be a topological space with the structure of a cell complex whose dimension is finite. We can define the dimension $\dim X$ uniquely, which is of course a non-negative integer. A topological manifold is well-known to be homeomorphic to some CW complex. A smooth manifold is well-known to be a polyhedron canonically, which is a so-called PL manifold. It is also well-known that a topological space homeomorphic to a polyhedron whose dimension is at most $2$ has the structure of a polyhedron uniquely. If we restrict this to the class of topological manifolds, then the condition is weaken to be that the dimension is at most $3$. Consult [22] for example.
@@
Spaces of smooth maps between given two manifolds are topologized with so-called $C^{\infty}$ Whitney topology. In short two smooth maps are mutually close if their differentials of arbitrary degrees are close. $C^{r}$ Whitney topologies are also considered as weaker topologies. They respect differentials of degrees at most $r$: $r$ is a non-negative integer of course. The $C^{r_{1}}$ Whitney topologies are weaker than the $C^{r_{2}}$ Whitney topologies for $0\leq r_{1}<r_{2}$: $r_{1}$ is a non-negative integer and $r_{2}$ is a positive integer or $\infty$. Such topologies are important in singularity theory and differential topology. See [8] for systematic exposition. Diffeomorphisms are defined as smooth maps being also homeomorphisms. The diffeomorphism group of a smooth manifold means the group of all diffeomorphisms from the manifold onto itself. It is also topologized with the $C^{\infty}$ Whitney topology as before of course. A smooth bundle means a bundle whose fiber is a smooth manifold and whose structure group is (a subgroup of) the diffeomorphism group. A linear bundle means a smooth bundle whose fiber is a Euclidean space, a unit disk, or a unit sphere and whose structure group consists of linear isomorphisms, defined naturally.
We present a theorem on fundamental structures of special generic maps. Using the notation, we also present our main result. Hereafter, for a manifold $X$, $\partial X$ denote the boundary of it.
@@
Let $m\geq n\geq 1$ be integers. Suppose that a smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n}}:\bar{N}\rightarrow{\mathbb{R}}^{n}$ from an $n$-dimensional compact and connected manifold $\bar{N}$ into ${\mathbb{R}}^{n}$ exists. Then we have a pair $(f:M\rightarrow{\mathbb{R}}^{n},f_{0}:M\rightarrow{\mathbb{R}}^{n})$ of maps on some $m$-dimensional closed and connected manifold $M$ which is also a non-singular real algebraic manifold and the following properties.
@@
    $f_{0}$ is a smooth real algebraic map.
@@
    There exists a pair $(\Phi,\phi)$ of homeomorphisms enjoying $\Phi\circ f_{0}\circ\phi=f$.
@@
Let $m\geq n\geq 1$ be integers. Suppose that a smooth immersion ${\bar{f}}_{N}:\bar{N}\rightarrow N$ from an $n$-dimensional compact manifold $\bar{N}$ into an $n$-dimensional smooth manifold $N$ with no boundary and the following additional data (objects) are given.
  • A small collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ and a smooth bundle ${\bar{M}}_{\rm Interior}={\bar{M}}_{\rm I}$ over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ whose fiber is the unit sphere $S^{m-n}\subset\partial D^{m-n+1}$. We call this bundle an interior unit sphere bundle for ${\bar{f}}_{N}$
@@
  • We consider the smooth bundle $\partial{\bar{M}}_{\rm I}$ as the restriction to the boundary $\partial(\bar{N}-{\rm Int}\ N(\partial\bar{N}))$. We also consider the subbundle $\partial{\bar{M}}_{\rm B}$ whose fiber is $S^{m-n}=\partial D^{m-n+1}$. We consider the identification between their base spaces in the natural way. More precisely, we can regard the collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ as the product $\partial\bar{N}\times[0,1]$. Furthermore, points $x\in\partial\bar{N}$ are identified with $(x,0)\in\partial\bar{N}\times\{0\}$ and $\partial N(\partial\bar{N})-\partial\bar{N}$ is regarded as the subspace of $\bar{N}-{\rm Int}\ N(\partial\bar{N})\subset\bar{N}$. The last object is a bundle isomorphism between the smooth bundle $\partial{\bar{M}}_{\rm I}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$.
We call the triple of an interior unit sphere bundle for ${\bar{f}}_{N}$, a boundary linear bundle for ${\bar{f}}_{N}$ and this bundle isomorphism a triple for a special generic map and ${\bar{f}}_{N}$ or an SGM-triple for ${\bar{f}}_{N}$. We have an $m$-dimensional closed manifold $M$ and a special generic map $f:M\rightarrow N$ such that $\bar{f}$ in Proposition 1 and ${\bar{f}}_{N}$ agree. Furthermore, we have this enjoying the following properties where we can abuse the notation in Proposition 1 and do.
@@
    We also consider the restriction of the interior unit sphere bundle ${\bar{M}}_{\rm I}$ to the boundary and the subbundle of the linear bundle whose fiber is $\partial D^{m-n+1}$ in Proposition 1. We consider their natural identification. This coincides with the bundle isomorphism in the SGM-triple for ${\bar{f}}_{N}$.
@@
In Proposition 2, if the interior unit sphere bundle and the boundary linear bundle are trivial and the bundle isomorphism in the SGM-triple for ${\bar{f}}_{N}$ is given as the product map of the identity map (identification) between the base spaces and the identity map on the fiber $S^{m-1}=\partial D^{m-n+1}$, then it is called a product-organized special generic map.
[15] has first introduced such a class of special generic maps. We first name the class in this way first in [12]. We do not need to understand this notion or [12, 15] well. However, we note that canonical projections of unit spheres are of such a class. We consider a manifold represented as a connected sum $\sharp(S^{k_{j}}\times S^{m-k_{j}})$ of finitely many manifolds $S^{k_{j}}\times S^{m-k_{j}}$ where $m\geq n\geq 2$ and $11\leq k_{j}\leq n-$ 1 are integers. We choose the connected sum in the smooth category. We can have a product-organized special generic map there into ${\mathbb{R}}^{n}$ from a natural embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ whose image is diffeomorphic to the boundary connected sum $\natural(D^{k_{j}}\times S^{n-k_{j}})$ of finitely many manifolds $D^{k_{j}}\times S^{n-k_{j}}$ taken in the smooth category. This is another simplest example.
We add the following as another main result.
@@
    $M$ is also represented as the zero set of some real polynomial.
@@
We prove our main result. STEPs 1–3 are main ingredients.
@@
is a smooth closed manifold, non-singular real algebraic manifold of dimension $m$ and the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$. The map $f_{0}$ is obtained as the composition of the canonical embedding of $M\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
@@
The resulting map is also special generic by the construction and product organized. $M$ is also seen to be ”diffeomorphic” to $M_{0}$. By considering the isotopy in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}$ enjoying all desired properties in Main Theorems.
@@
In [10, 11, 14], we consider cases where polynomials are given beforehand. In short, only STEP 2 and STEP 3 are considered. These studies also consider explicit cases. For example, polynomials are of degree $2$ and for spheres. These studies are regarded as studies finding nice classes of real algebraic maps including the canonical projections of unit spheres as simplest cases.
@@
The auhtor would like to thank Takahiro Yamamoto for asking some questions on the talk by the author, given on 2023/11/27, and related informal discussions at Kyushu University held on 2023/12/11, after the conference. He knows singularity theory and applications to geometry widely. Based on his knowledge and idea, he has first commented that projective spaces admit no special generic maps in considerable cases and that it is interesting: this is on the study [12] of the author. He also says that it is also important to consider so-called general fold maps and explicit classes for understanding projective spaces in geometric ways: fold maps are defined naturally generalizing the form of the last ($n$-th) component in the definition of a special generic map (see related pioneering studies [33, 35], studies [5, 6] on existence, [26], a pioneering studies on algebraic topology and differential topology of manifolds admitting fold maps, and [8] for related fundamental and advanced theory from singularity theory for example).

### 09 Jan 2024 v2

# A note on real algebraic maps which are topologically special generic maps

@@ Abstract.
Morse functions with exactly two singular points on homotopy spheres and canonical projections of spheres are generalized as special generic maps. A special generic map is, roughly, a smooth map represented as the composition of a smooth surjection onto a manifold whose preimages are diffeomorphic to a unit sphere in the interior of the manifold and single point sets on the boundary with a smooth immersion of codimension 00.
@@
In theory of Morse functions, related singularity theory and applications to geometry of manifolds, Morse functions with exactly two singular points on spheres (in Reeb’s theorem) are of most fundamental functions and maps. Such maps are generalized to special generic maps: they are smooth maps locally represented as projections or singular points of the form $(x_{1},\cdots x_{m})\mapsto(x_{1},\cdots x_{n-1},{\Sigma}_{j=1}^{m-n+1}{x_{j}}^{2})$. [4, 7, 28] are related pioneering studies. For Morse functions, see [26] for a pioneering study and [21] for systematic exposition on fundamental notions and arguments.
@@
For example, [19, 25, 33] shows explicit cases and some fundamental properties of real algebraic functions in the theory of Lie groups and symmetric spaces. (Unit) spheres are simplest symmetric spaces. We are interested in another generalizations of projections of spheres: constructing real algebraic maps topologically special generic maps. The author has succeeded in such construction in [10], followed by [14] by the author himself. We present results of a similar type which are also new.
@@
Let $X$ be a topological space with the structure of a cell complex whose dimension is finite. We can define the dimension $\dim X$ uniquely, which is of course a non-negative integer. A topological manifold is well-known to be homeomorphic to some CW complex. A smooth manifold is well-known to be a polyhedron canonically, which is a so-called PL manifold. It is also well-known that a topological space homeomorphic to a polyhedron whose dimension is at most $2$ has the structure of a polyhedron uniquely. If we restrict this to the class of topological manifolds, then the condition is weaken to be that the dimension is at most $3$. Consult [23] for example. Hereafter, for a manifold $X$, $\partial X$ denote the boundary of it.
@@
Spaces of smooth maps between given two manifolds are topologized with so-called $C^{\infty}$ Whitney topologies. In short two smooth maps are mutually close if their differentials of arbitrary degrees are close. $C^{r}$ Whitney topologies are also considered as weaker topologies. They respect differentials of degrees at most $r$: $r$ is a non-negative integer of course. The $C^{r_{1}}$ Whitney topologies are weaker than the $C^{r_{2}}$ Whitney topologies for $0\leq r_{1}<r_{2}$: $r_{1}$ is a non-negative integer and $r_{2}$ is a positive integer or $\infty$. Such topologies are important in singularity theory and differential topology. We omit rigorous exposition. See [8] for systematic exposition. Diffeomorphisms are defined as smooth maps being also homeomorphisms. The diffeomorphism group of a smooth manifold means the group of all diffeomorphisms from the manifold onto itself. It is also topologized with the $C^{\infty}$ Whitney topology as before of course. A smooth bundle means a bundle whose fiber is a smooth manifold and whose structure group is (a subgroup of) the diffeomorphism group. A linear bundle means a smooth bundle whose fiber is a Euclidean space, a unit disk, or a unit sphere and whose structure group consists of linear isomorphisms, defined naturally.
We present a theorem on fundamental structures of special generic maps. Using the notation, we also present our main result.
@@
Let $m\geq n\geq 1$ be integers. Suppose that a smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n}}:\bar{N}\rightarrow{\mathbb{R}}^{n}$ from an $n$-dimensional compact and connected manifold $\bar{N}$ into ${\mathbb{R}}^{n}$ exists. Then we have a pair $(f:M\rightarrow{\mathbb{R}}^{n},f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n})$ of maps on two suitably obtained mutually homeomorphic $m$-dimensional closed and connected manifolds $M$ and $M_{0}$ with the following properties.
@@
    $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}$.
@@
    There exists a pair $(\Phi;M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms enjoying $\phi\circ f_{0}\circ\Phi=f$.
@@
Let $m\geq n\geq 1$ be integers. Suppose that a smooth immersion ${\bar{f}}_{N}:\bar{N}\rightarrow N$ from an $n$-dimensional compact manifold $\bar{N}$ into an $n$-dimensional smooth manifold $N$ with no boundary and the following additional data (objects) are given. Hereafter we can abuse the notation in Proposition 1 and do.
  • A small collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ and a smooth bundle ${\bar{M}}_{\rm Interior}={\bar{M}}_{\rm I}$ over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ whose fiber is the unit sphere $S^{m-n}\subset\partial D^{m-n+1}$. We call this bundle an interior unit sphere bundle for ${\bar{f}}_{N}$.
@@
  • We consider the smooth bundle $\partial{\bar{M}}_{\rm I}$ as the restriction to the boundary $\partial(\bar{N}-{\rm Int}\ N(\partial\bar{N}))$. We also consider the subbundle $\partial{\bar{M}}_{\rm B}$ whose fiber is $S^{m-n}=\partial D^{m-n+1}$. We consider the identification between their base spaces in the natural way. More precisely, we can regard the collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ as the product $\partial\bar{N}\times[0,1]$. Furthermore, points $x\in\partial\bar{N}$ are identified with $(x,0)\in\partial\bar{N}\times\{0\}$ and $\partial N(\partial\bar{N})-\partial\bar{N}$ is regarded as the subspace of $\bar{N}-{\rm Int}\ N(\partial\bar{N})\subset\bar{N}$. The last object is a bundle isomorphism between the smooth bundle $\partial{\bar{M}}_{\rm I}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$. We call it an identification bundle isomorphism for ${\bar{f}}_{N}$.
We have an $m$-dimensional closed manifold $M$ and a special generic map $f:M\rightarrow N$ such that $\bar{f}$ in Proposition 1 and ${\bar{f}}_{N}$ agree. Furthermore, we have this enjoying the following properties.
@@
    We also consider the restriction of bundle ${\bar{M}}_{\rm I}$ in (1) here to the boundary and the subbundle of the linear bundle whose fiber is $\partial D^{m-n+1}$ in (2) here. We consider their natural identification. This coincides with the identification bundle isomorphism for ${\bar{f}}_{N}$.
@@
In Proposition 2, if the interior unit sphere bundle and the boundary linear bundle for ${\bar{f}}_{N}$ are trivial and the identification bundle isomorphism for ${\bar{f}}_{N}$ is given as the product map of the identity map (identification) between the base spaces and the identity map on the fiber $S^{m-1}=\partial D^{m-n+1}$, then it is called a product-organized special generic map.
[15] has first introduced such a class of special generic maps. We first name the class in this way first in [12]. Of course in our present paper, we do not need to understand this notion or [12, 15] well and we do not assume their arguments or results.
However, we note that canonical projections of unit spheres are of such a class. We consider a manifold represented as a connected sum $\sharp(S^{k_{j}}\times S^{m-k_{j}})$ of finitely many manifolds $S^{k_{j}}\times S^{m-k_{j}}$ where $m\geq n\geq 2$ and $11\leq k_{j}\leq n-$ 1 are integers. We choose the connected sum in the smooth category. We can have a product-organized special generic map there into ${\mathbb{R}}^{n}$ from a natural embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ whose image is diffeomorphic to the boundary connected sum $\natural(D^{k_{j}}\times S^{n-k_{j}})$ of finitely many manifolds $D^{k_{j}}\times S^{n-k_{j}}$ taken in the smooth category. This gives another simplest example.
We add the following as additional main results.
@@
    $M_{0}$ is also represented as the zero set of some real polynomial.
Main Theorem 3.
In Main Theorem 2, we can have $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ with the following properties.
  1. (1)
    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” and have a smooth bundle whose projection is defined as the composition of the restriction of $q_{f_{0}}$ to the preimage of $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and whose structure group is reduced to a group consisting of linear transformations of $D^{m-n+1}$: in fact a manifold homeomorphic to $D^{k}$ is known to be diffeomorphic to $D^{k}$ and this is a linear bundle as defined except for $m-n+1=k\neq 4$.
  2. (2)
    Under the structure group, the bundle of (1) here is in fact trivial.
  3. (3)
    $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ diffeomorphic to a manifold obtained in the following steps.
    1. (a)
      Prepare the product $\bar{N}\times D^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$ for a manifold ${\Delta}^{m-n+2}$ homeomorphic to $D^{m-n+2}$. Note that ${\Delta}^{k}$ is diffeomorphic to $D^{k}$ for $k\neq 4$.
    2. (b)
      Attach the previous two manifold by a bundle isomorphism between the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}$ over $\partial\bar{N}$ and the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}\subset(\partial\bar{N})\times\partial{\Delta}^{m-n+2}$ over $\partial\bar{N}$ represented as the product of the identity maps on $\partial\bar{N}$ and $D^{m-n+1}$: here $D^{m-n+1}$ is smoothly embedded in $\partial{\Delta}^{m-n+2}$.
    3. (c)
      Eliminate the corner.
    In fact, except the case where ${\Delta}^{k}$ is not diffeomorphic to $D^{k}$, $M$ and $M_{0}$ are diffeomorphic.
Remember that ”${\Delta}^{k}$ not being diffeomorphic to $D^{k}$” is still undiscovered.
@@
We prove our main result. First we prove Main Theorems 1 and 2. STEPs 1–3 are main ingredients.
@@
is a smooth closed manifold, non-singular real algebraic manifold of dimension $m$ and the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$. The map $f_{0}$ is obtained as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
@@
The resulting map is also special generic by the construction and product organized. By considering the isotopy in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}$ enjoying all desired properties in Main Theorems 1 and 2. This completes the proof of Main Theorems 1 and 2.
We show Main Theorem 3. We can prove this by applying (revised versions of) known arguments previously presented in related studies and our present study.
We first show Main Theorem 3 (1). First the bundle we discuss here is a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this). We apply an argument used to show Proposition 1. More precisely, Proposition 2.1 of [28] is a corresponding result. It also refers to [22] there. The structure group is regarded as a group consisting of diffeomorphisms preserving the distance between each point and the origin $0\in{\Delta}^{m-n+1}$ where we can define the origin 00 for ${\Delta}^{m-n+1}$ not diffeomorphic to $D^{m-n+1}$. We explain about this. For an element ${\phi}_{{\Delta}^{m-n+1}}$ of such a group, we consider $\frac{1}{t}{\phi}_{{\Delta}^{m-n+1}}(tx)$ for $x\in{\Delta}^{m-n+1}$ and $0<t\leq 1$ where $x$ and $tx$ are canonically regarded as vectors. Here $||\frac{1}{t}{\phi}_{{\Delta}^{m-n+1}}(tx)||=||x||$. We consider the limit at $t=0$. For a fixed $x$, we have the value of the differential of ${\phi}_{{\Delta}^{m-n+1}}$ at $x$, seen naturally as a tangent vector at the origin of ${\Delta}^{m-n+1}$. Thus we naturally have an element regarded as a linear transformation of ”the unit disk $D^{m-n+1}$”. This means that the structure group reduces to the structure group of a linear bundle whose fiber is $D^{m-n+1}$. This completes the proof of Main Theorem 3 (1).
We see that this bundle is trivial for the structure group.
$W_{0}$ is defined as
$\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}\geq 0\}$
and this can be also defined as
$\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}\geq 0\}$.
We use the function $g$ before. We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{g(t)}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule where we consider some elements as vectors.
  • If $x\in{\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
From this we can see that $W_{0}-S(f_{0})$ is regarded as the total space of a trivial smooth bundle over ${\rm Int}\ {\bar{N}}_{0}$ whose fiber is diffeomorphic to $D^{m-n+1}$. By the structure of the map around $S(f_{0})$ and $\partial{\bar{N}}_{0}$, we can see that the previous bundle is trivial for the structure group. This completes the proof of Main Theorem 3 (2).
By our arguments, we can also check Main Theorem 3 (3) easily. A result close to this is also presented in Theorem 1 (3) of [14]. ”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$ there. The topology of the ($m+1$)-dimensional manifold is not so precisely discussed there: only exposition on ”homotopy” is presented there.
@@
In [10, 11, 14], we consider cases where polynomials are given beforehand. In short, only STEP 2 is considered and related short exposition on STEP 3 is presented where the latter presentation is a kind of short remark and this does not affect main ingredients of [10, 11, 14]. STEP 3 explicitly shows that $f_{0}$ is ”topologically special generic” first. These studies also consider explicit cases. For example, polynomials are of degree $2$ and for spheres. These studies are regarded as studies finding nice classes of real algebraic maps including the canonical projections of unit spheres as simplest cases.
@@
The following is a remark on Main Theorem 3.
Remark 4.
As another method, we can also show Main Theorem 3 in the case $m=n$ by respecting the well-known facts that the diffeomorphism group of $D^{1}$ has the homotopy type of the two-point discrete set and that each point of the two-point set corresponds to orientation preserving diffeomorphisms and orientation reversing ones, respectively. In this case, $M$ and $M_{0}$ are diffeomorphic to the so-called double of $\overline{N_{0}}$.
@@
The author would like to thank Takahiro Yamamoto for asking some questions on the talk by the author, given on 2023/11/27, and related informal discussions at Kyushu University held on 2023/12/11, after the conference. He knows singularity theory and applications to geometry widely. Based on his knowledge and idea, he has first commented that projective spaces admit no special generic maps in considerable cases and that it is interesting: this is on the study [12] of the author. He also says that it is also important to consider so-called general fold maps and explicit classes for understanding projective spaces in geometric ways: fold maps are defined naturally generalizing the form of the last ($n$-th) component in the definition of a special generic map (see related pioneering studies [34, 36], studies [5, 6] on existence, [27], a pioneering studies on algebraic topology and differential topology of manifolds admitting fold maps, and [8] for related fundamental and advanced theory from singularity theory for example).